# Inferring State Machine from the Protocol Implementation via Large Language Model

### v3

@@ Abstract
State machines play a pivotal role in augmenting the efficacy of protocol analyzing to unveil more vulnerabilities. However, inferring state machines from network protocol implementations presents significant challenges, mainly because of the complicated code syntax and semantics. Traditional methods based on dynamic analysis often overlook crucial state transitions due to limited coverage, while static analysis suffers from path explosion facing to protocol implementations. To address these limitations, we propose an innovative state machine inference approach powered by Large Language Models (LLMs) named ProtocolGPT. Utilizing retrieval augmented generation technology, this method augments pre-trained model with specific knowledge drawn from protocol implementations. Through targeted prompt engineering, we systematically identify and infer the underlying state machines. Our evaluation across six protocol implementations demonstrates the method’s high efficacy, achieving precision exceeding 90% and successfully delineating differences on state machines among various implementations of the same protocol. Integrating our approach with protocol fuzzing significantly improves fuzzers by more than 20% in terms of coverage and detects two zero-day vulnerabilities compared to the baseline. Our proposed method represents a major advancement in accurate state machine inference and highlights the substantial potential of LLMs in enhancing network protocol security analysis.
@@ I Introduction
Finite State Machines (FSMs) are essential components of protocols, comprising various states and the transitions between these states based on specific conditions or inputs, which guide the implementation of protocols [1]. FSM inference serves as a fundamental cornerstone in a wide range of applications, including vulnerability detection, software engineering, and network protocol analysis [2] while improper protocol implementations may lead to serious security issues [3] and most of the vulnerabilities in network protocols are intricately linked to these states [4]. Therefore, accurately inferring protocol state machines is critical for understanding protocol functionalities [5], protocol verification [6], and unveiling vulnerabilities [7].
@@
Most protocol implementations adhere to the specifications outlined in the network protocol RFCs, which define communication functions, input formats, and state machines [8]. Recently, some researchers have employed natural language processing (NLP) techniques to infer state machines from these specifications [9][8]. Due to the inherent subjectivity in the definitions within RFCs, different developers may interpret them in varying ways. As a result, protocol implementations often include custom elements, leading to discrepancies between the FSMs in protocol implementations and those defined in the RFCs. Even different implementations of the same protocol can exhibit variations considerably in their FSMs, which are the most often hotspots for vulnerabilities. Therefore, the state machines extracted from specific protocol implementations instead of RFCs are more precise and important for protocol security analysis, and can unleash more potential for protocol security analysis.
@@
Research on inferring state machines from protocol implementations can be categorized into two principal types: static analysis [10] and dynamic analysis [11] [12] [13] [14] [15] [16]. Static analysis methods, such as control flow analysis, and data flow analysis, infer state machines by deeply understanding the code structure, eliminating the need for program execution. While static analysis provides deep insights without running the code, it struggles with complicated code structures and uncertain dynamic behavior contexts, often resulting in the path explosion problem. Conversely, dynamic analysis extracts state machines by tracing program behavior and monitoring runtime data during execution. Although dynamic analysis is beneficial for capturing real-time program behavior, it may overlook critical state transitions due to limited coverage, and its effectiveness significantly depends on the quality and representativeness of test cases. These limitations underscore the need for a scalable and comprehensive method to infer state machines from protocol implementations.
Recent studies have demonstrated that LLM possess extensive potential for applications in program analysis. By generating both natural and programming languages, they can assist developers with tasks such as code writing, code repair, vulnerability detection, code interpretation, and documentation generation [17] [18] [19] [20] [21]. Moreover, LLMs have the capability to guide protocol fuzzing [22]. However, these usages has been limited to generating protocol inputs, without the ability to infer the protocol finite state machine. And we have found that LLMs can infer protocol states and state transitions within program code to a certain extent in several cases. Based on the observations above, we posit that LLMs possess the potential to deduce state machines from protocol implementations. In this paper, we primarily investigate FSM in protocol implementations, utilize LLMs to identify protocol states, message types, and transition relationships from the source code. Limited by context size limitations, sufficient information will not be provided to LLMs resulting the inefficient for LLMs to analyze complicated codebases.
@@
Moreover, as the volume and complexity of the code increase, the accuracy of LLMs correspondingly decreases. To address this dilemma, we propose a method for inferring FSMs from protocol implementations using LLM augmented with source code. First, we filter out FSM-related code from large and complicated protocol implementations to avoid unnecessary analysis. Subsequently, the FSM-related code is processed through an embedding model and converted into a vector store which is utilized as augmenting the pre-trained LLM. Finally, we guide the augmented LLM to infer protocol states, message types, and state transitions using chain-of-thought techniques [23].
@@
Our proposed approach aims to overcome several prevailing obstacles in inferring protocol FSMs, offering distinct advantages over traditional techniques such as dynamic program analysis, which typically depend on network traffic samples, and static analysis, which is generally limited by the situation mentioned above. Moreover, our method effectively infers the authentic FSMs from protocol implementations, enabling the identification of discrepancies between the actual protocol implementations and their respective RFC specifications.
To rigorously examine the proficiency of LLM in inferring state machines from protocol implementations, we meticulously infer the state machines of six distinct protocols: IKEv2, TLS1.3, TLS1.2, BGP, RTSP, and L2TP. The empirical findings from our study reveal that the state machines inferred via our methodology exhibit an average precision rate exceeding 90%. Notably, within these deduced state machines, we discern variances across disparate implementations of identical protocols. Furthermore, we apply the inferred FSMs to augment the efficacy of the protocol fuzzing tool AFLNet, thereby facilitating enhance code coverage. In essence, our contributions are threefold:
  • We proposed a method to augment LLM with protocol implementation. This approach addresses limitations encountered by LLMs while analyzing source code.
@@
  • We pioneered a new approach, which was not proposed before, to infer protocol FSMs using augmented LLMs. This novel method has the capability to infer protocol states, message types, and state transitions from protocol implementations.
  • We implemented our prototype, named  ProtocolGPT, which can automatically infer FSMs from protocol implementations. We designed and conducted experiments demonstrating that the augmented LLMs achieve precision exceeding 90% in state machine inference. Through the application of  ProtocolGPT, we discovered discrepancies within FSMs between different implementations of the same protocol. Furthermore, AFLNet enhanced by  ProtocolGPT increases the coverage of statements by 22. 23% on average compared to baselines. We discovered two 0-day vulnerabilities and 18 previously known vulnerabilities.
@@ II Background and Related Work
In this section, we introduce protocol state machine inference techniques and large language models. We start by presenting several protocol state machine inference techniques, highlighting their respective advantages and limitations. Subsequently, we provide some background on large language models.
II-A Protocol State Machine Inference
Figure 1: The state machine of the TLS 1.3 protocol server defined in RFC 8446.
The protocol state machine is a model used to describe the logic and behavior of interaction between peers. The essential components of a protocol state machine include the states of peers, inputs, and state transitions. The set of protocol states encompasses all possible states that can be reached by peers during the interactions. Input refer to all messages that can be sent or received by the server and client. Each protocol specification defines its unique input format, and both parties must strictly adhere to this format during interactions. The reception or transmission of a message at the communication terminal may lead to a state transition. The state machine of TLS 1.3 server is shown in Figure 1, highlighting states like START, RECVD_CH, and NEGOTIATED, and detailing the exchange of messages such as ClientHello, ServerHello, and Certificate. This interaction initiates in the START state and progresses through a series of exchanges to the CONNECTED state, establishing a secure connection between the server and the client.
Previous efforts to infer protocol state machine include dynamic analysis [11] [12] [13] [14] [15] [16], static analysis [10], and natural language processing  [9] [24] [25]. Dynamic program analysis investigates the data flow or control flow during message parsing, leveraging previously captured network traffic and utilizing technologies such as statistical method and active learning to deduce protocol formats. Conversely, static analysis engages in a direct examination of source code, employing symbolic execution and data flow analysis to infer protocol message formats. NLP, meanwhile, delves into the elucidation of transition relationships between protocol states using natural language inputs, such as RFCs and technical documentation.
II-B Large Language Model
@@
Large language models are a cutting-edge advancement in artificial intelligence, rooted in deep learning and natural language processing. Capable of performing tasks like text generating [26] [27] [28], translating [29] [30] [31], and summarizing text [32] [33], these models undergo a process of pre-training and fine-tuning to grasp the intricate structure and semantics of language. Recently, LLMs have excelled in software engineering, including code generating [34] [35], code repairing [36] [37] [38], and vulnerability detection [39] [18]. Almost all applications of LLMs in software engineering are based on prompt engineering and fine-tuning. With massive train data contained network protocol and source code, LLMs have the potential to deduce state machines from protocol implementations, showcasing their pivotal role in enhancing protocol security and efficiency.
LLMs demonstrate exceptional potential in program analysis. Nevertheless, they encounter specific challenges in conducting program analysis [40]. Firstly, the output generated by LLMs may occasionally manifest randomness, leading to the creation of non-existent facts. Secondly, in the face of intricate analysis tasks, LLMs are incapable of rendering efficient completions in a singular effort, necessitating incremental guidance via human-formulated prompts. Moreover, in scenarios involving complicated and advanced programs, the precision of LLMs falls short when compared to some conventional analysis methodologies. This discrepancy can largely be ascribed to the constrained scope of LLMs’ context windows, which hampers their ability to thoroughly examine highly complicated and voluminous codebases in a singular analytical session.
@@ III Motivation and Case Study
In this section, we use several cases as motivations to demonstrate the reasons and challenges of inferring state machines from the protocol implementations via LLM.
III-A Inconsistencies between Protocol Implements and RFCs
To investigate the inconsistencies between protocol specifications and implementations, we conduct a manual analysis of four implementations of the IKEv2 protocol, in conjunction with RFC 7296. Table I provides a summary of the number of states and state transitions for each implementation. To ensure the reliability of our results, two domain experts thoroughly examined both the RFC and the protocol implementations.
As illustrated in the table, there are marked discrepancies between the state machines described in the RFC and those observed in the protocol implementations. Several factors contribute to these discrepancies. Firstly, protocol specifications are typically documented in written form, leading to varying interpretations and understandings among different developers or teams. This variability can result in significant differences between the FSMs of the protocol implementations and those defined in the RFC. Additionally, the official RFC guidelines for communication mechanisms include several open-ended definitions, allowing developers to design protocol implementations tailored to their specific application scenarios and functional requirements. Furthermore, developers may introduce custom features or extensions to the protocol implementations to address particular needs, which may not be supported by other implementations, thus leading to inconsistencies.
Consequently, there are inherent inconsistencies between the state machines of RFC and implementations. The state machines inferred from implementations, as opposed to those strictly defined in the RFC, can provide better insights for analyzing software behavior. However, as highlighted in Section II, inferring state machines from implementations remains a complicated challenge.
Finding 1: Discrepancies between specifications and protocol implementations are ubiquitous across various systems.
III-B Inferring protocol state machine via LLM
LLMPF [22] leverages the syntactic information generated by the LLM to enrich the seeds corpus used for fuzzing. Rather than utilizing the LLM augmented with specific knowledge, LLMPF directly employs a pre-trained model. To the best of our knowledge, augmented models can deal with more specific problems than pre-trained models. We thus compare the pre-trained model with the augmented model by employing a simple state machine inference case. For the pre-trained LLM, we adopt an approach akin to LLMPF, employing prompt engineering to incrementally guide the model in generating protocol states and transitions. In contrast, we utilize a pre-trained large language model augmented with source code to infer the state transitions. Initially, the pre-trained model is enhanced with the protocol implementation, followed by the application of the chain-of-thought technique to guide the augmented LLM in inferring the state machine. The methodology for state machine inference using the augmented LLM will be elaborated in Section IV. For the IKE protocol implementation, we selected strongSwan[41], a widely used solution in the industry.
@@ III-B Inferring protocol state machine via LLM
TABLE II: The comparison between state machines inferred by pre-trained model and augmented model.
Models States Transitions Precision
Pre-trained LLM 4 12 43.48%
Augmented LLM 8 20 95.00%
Table II presents a comparison of the state machine inferred by the pre-trained model and the augmented model. The detailed experimental methodology is discussed in Section V. The result shows that the augmented model outperforms the pre-trained model by 51.52% in terms of precision. The enhanced LLM demonstrates the capability to extract reliable state machine information from the source code. While the pre-trained LLM can offer some insights related to FSMs, its precision is limited to 43.48%. This discrepancy may be attributed to the hallucinations exhibited by LLMs when applied to domain-specific tasks.
Finding 2: Compared to the pre-trained model, the LLM augmented with source code exhibits a significantly enhanced capacity to generate a more reliable state machine.
III-C Context Window of LLM
TABLE III: Detail information of benchmark protocols.
@@ III-C Context Window of LLM
The performance of LLM is heavily dependent on the context information provided [50]. However, the length of context is limited by the input capacity of model. Moreover, the performance of model tends to degrade with an increase in input length. This limitation applies to both closed-source and open-source models. For example, OpenAI’s GPT-4 model has a context window of only 8192 tokens. While this input length suffices for typical NLP tasks, it is inadequate for program analysis. As illustrated in Table III, the number of tokens required for the implementations of protocols far exceeds the context window of the GPT-4 model.
When inferring state machines directly from implementation using pre-trained model, the complexity of the code and the sheer volume of files involved can hinder the model from producing accurate results. Protocol implementations typically comprise components such as state machines, configurations, tests, and encryption mechanisms. However, the code related to the state machine is often concentrated in specific files. For instance, the message types are usually defined in a single file. Identifying these specific files can narrow the scope of the code to be analyzed. Conversely, analyzing code related to configurations, encryption, authentication, and testing does not yield relevant information about the state machine itself.
Finding 3: The limitation of context window renders inferring state machine via LLM a challenging endeavor.

### v4

# Unleashing the Power of LLM to Infer State Machine from the Protocol Implementation

@@ Abstract
State machines are essential for enhancing protocol analysis to identify vulnerabilities. However, inferring state machines from network protocol implementations is challenging due to complex code syntax and semantics. Traditional dynamic analysis methods often miss critical state transitions due to limited coverage, while static analysis faces path explosion issues. To overcome these challenges, we introduce a novel state machine inference approach utilizing Large Language Models (LLMs), named ProtocolGPT. This method employs retrieval augmented generation technology to enhance a pre-trained model with specific knowledge from protocol implementations. Through effective prompt engineering, we accurately identify and infer state machines. To the best of our knowledge, our approach represents the first state machine inference that leverages the source code of protocol implementations. Our evaluation of six protocol implementations shows that our method achieves a precision of over 90%, outperforming the baselines by more than 30%. Furthermore, integrating our approach with protocol fuzzing improves coverage by more than 20% and uncovers two 0-day vulnerabilities compared to baseline methods.
@@ I Introduction
Finite State Machine (FSM) is a core component of protocols, comprising multiple states and transitions based on specific inputs. FSM inference is the fundamental building block of various application domains, including vulnerability detection, software engineering, and network protocol analysis. Improper protocol implements can lead to serious security issues, potentially leading to vulnerabilities [1]. So accurately inferring protocol FSMs is crucial for understanding functionality, verification, and finding vulnerabilities[2][3][4].
@@
Most protocol implementations follow the specifications of network protocol (RFC, i.e., Request for Comments), which define the standard communication functions, input formats, and FSMs [5]. In recent years, researchers have used natural language processing (NLP) techniques to infer FSMs from these specifications [6][5]. However, different developers may have discrepant interpretations on the same RFC, resulting in discrepancies between the implemented FSM and the RFC. Even different implementations of the same protocol may have significantly discrepant FSMs, and these discrepancies are often hotspots for vulnerabilities. Therefore, FSMs extracted from specific protocol implementations are more accurate than those in RFCs, more important for protocol security analysis, and can unlock greater potential.
@@
Research on inferring FSMs from protocol implementations can be categorized into two principal types: static analysis [7] and dynamic analysis [8] [9] [10] [11]. Static analysis approaches, such as control flow analysis, and data flow analysis, infer FSMs by deeply understanding the code structure, eliminating the need for program execution. While static analysis provides deep insights without running the code, it struggles with complicated code structures and uncertain dynamic behavior contexts, often resulting in the path explosion problem. Conversely, dynamic analysis extracts FSMs by tracing program behavior and monitoring runtime data during execution. Although dynamic analysis is beneficial for capturing real-time program behavior, it may overlook critical state transitions due to limited coverage, and its effectiveness significantly depends on the quality of test cases. These limitations underscore the need for a scalable and comprehensive method to infer FSMs from protocol implementations.
Recent studies have shown that LLM has extensive potential for applications in program analysis, which can assist developers with tasks such as programming, code repair, vulnerability detection [12] [13], even guiding protocol fuzzing [14]. However, applying LLM to infer protocol FSM from source code implementations still encounters several challenges: ❶ The protocol implementations are highly complex and with great code amount. The core components should be precisely paid attention for analysis, while the irrelavant parts should be excluded preventing from side effects. ❷ The tokens of implementations significantly exceed the input size of LLMs. Effectively leveraging the knowledge embedded within the repositories to enhance LLM remains a substantial challenge. ❸ The procedure for the FSM inference of the protocol is complicated, due to the diverse appearance of static and dynamic status, which cannot be accomplished within a step.
@@
To tackle these challenges, we propose a novel perspective on inferring FSMs from protocol implementations by augmenting LLMs with source code. Initially, we isolate FSM-related code from implementations to eliminate irrelevant analysis. The extracted code is then partitioned into appropriately sized segments. Then these segments are converted into a vector store, which augments the pre-trained LLM. Finally, we employ chain-of-thought (CoT) [15] and background-augmented prompting (BAP) [16] techniques to guide the augmented LLM in inferring protocol states, message types, and state transitions.
@@
Our proposed approach aims to overcome several prevailing obstacles in inferring protocol FSMs, offering distinct advantages over traditional techniques such as dynamic program analysis, which typically depends on network traffic samples, and static analysis, which is generally limited by the situation mentioned above. Moreover, our method effectively infers the authentic FSMs from protocol implementations, enabling the identification of discrepancies between the actual protocol implementations and their respective RFC specifications.
To evaluate our approach, we conducted inferences on the FSMs of six protocols (IKEv2, TLS1.3, TLS1.2, BGP, RTSP, and L2TP). The evaluation results demonstrate that ProtocolGPT achieves a over 30% improvement in precision and a over 55% enhancement in recall compared to the baselines. In the inferred FSMs, we observed discrepancies between implementations of the same protocol. Additionally, the fuzz tester enhanced by ProtocolGPT achieved over 20% higher code coverage and discovered two zero-day vulnerabilities. Our contribution can be summarized as follows:
  • We propose a novel method to enhance LLMs with protocol implementation knowledge, effectively addressing the limitations faced by LLMs in source code analysis.
@@
  • We firstly introduce an approach for inferring protocol FSMs via implementation with augmented LLMs, enabling the identification of protocol states, message types, and state transitions from protocol implementations.
  • We developed ProtocolGPT, achieving over 90% precision in FSM inference. This tool reveals discrepancies in FSMs across discrepant implementations of the same protocol. Additionally, when integrated with AFLNet, ProtocolGPT improves code coverage by over 20% and discover two 0-day vulnerabilities which are never detected by the others before.
@@ II Background and Related Work
In this section, we introduce inference techniques of the protocol’s state machine, from rule-based to learning-based.
II-A Rule-based Protocol State Machine Inference
The protocol state machine is a model used to describe the logic and behavior of interaction between peers. Conventional efforts to infer protocol state machine include static analysis [7] and dynamic analysis [8] [9] [10] [11].
Static analysis involves directly examining source code using formal methods such as symbolic execution or data flow analysis in protocol reverse engineering. For instance, Shi et al. use abstract interpretation to extract protocol formats from parsers within protocol implementations [7]. However, static analysis methods are limited in terms of generality and scalability, often facing state explosion issues when applied to complex programs. Furthermore, static analysis is not applicable when the source code is unavailable. In contrast, dynamic analysis methods do not require access to the source code. These methods examine data or control flow during message parsing, leveraging previously captured network traffic and applying techniques such as statistical methods and active learning to infer state machines. For instance, Kleber et al. proposed a strategy for classifying message types in binary protocols [10] using a similarity metric based on continuous value ranges. Ye et al. introduced a probabilistic approach based on joint distribution for reversing network protocols [9]. Nevertheless, all of these methods share the common limitation of limited coverage.
II-B Learning-based Protocol State Machine Inference
In recent years, machine learning has gained significant attention for its potential to automate and enhance the process of protocol state machine inference [6]  [17]  [18]. Unlike traditional methods, which often rely on manually crafted rules or heuristics, machine learning approaches leverage data-driven models to learn patterns from observed protocol behaviors. Pacheco et al. proposed a deep learning model based on a natural language data corpus to infer protocol state machines [6]. However, these data-driven approaches cannot be applied to protocol implementations in the same way as program analysis methods.
@@ II-B Learning-based Protocol State Machine Inference
Large language models are a cutting-edge advancement in artificial intelligence, underpinned by deep learning and natural language processing. Recently, LLMs have excelled in software engineering, including code generating  [19], code repairing  [20], and vulnerability detection [21]. Furthermore, Meng et al. leverage LLMs to enrich the seed corpus and overcome coverage plateau in protocol fuzzing [14]. Therefore, LLMs have the potential to deduce state machines from protocol implementations, showcasing their pivotal role in enhancing protocol security and efficiency.
@@ III Motivation and Case Study
Although LLMs are powerful, in this section, we use several cases as motivations to demonstrate the reasons and challenges of inferring FSM from the protocol implementations via LLM.
III-A Discrepancies between Protocol Implements
Discrepancies often arise between different protocol implementations from the same RFC, which creates great resistance for inferring FSM from source code. Opportunity and challenge coexist, which indicates that inferring from source code implements is more useful for real world analysis. To illustrate this issue, we conduct a manual analysis of four implementations of the IKEv2 protocol, comparing them with RFC. Table I provides a summary of the number of states and transitions. To ensure the reliability of our results, two domain experts thoroughly examined both the RFC and the protocol implementations.
@@ III-A Discrepancies between Protocol Implements
As illustrated in the table, there are marked discrepancies between the state machines defined in the RFC and those observed in the protocol implementations. Several factors contribute to these discrepancies. Firstly, protocol specifications are typically documented in written form, leading to varying interpretations among different developers. This variability can result in significant discrepancies between the protocol implementations and the RFC. Additionally, the official RFC guidelines for communication mechanisms include several open-ended definitions, allowing developers to design applications tailored to their specific scenarios and functional requirements. Furthermore, developers may introduce custom features to meet specific needs, which may not be supported by other implementations. Consequently, there are inherent discrepancies between the RFCs and implementations. The FSMs inferred from implementations, as opposed to those defined in the RFC, can provide better insights for analyzing software behavior. However, as highlighted in Section II, inferring FSMs from implementations remains a challenge.
Finding 1: Due to discrepancies between protocol implementations and RFCs, inferring the FSM through source code implementation is more effective for software analysis than RFC-based derivation.
III-B Limitation of LLM’s Input Length
LLM is powerful, but LLM heavily relies on the context information [27]. Even leveraging the source code implementation to infer FSM is better, but the limted length of context of the current LLMs’ remains a great challenge. For example, GPT-4 has only a context window of 8,192 tokens. This input length suffices for majority of NLP tasks, not enough for program analysis. As illustrated in Table II, the tokens of protocol implementations far exceeds the context window of the GPT-4.
TABLE II: Detail information of benchmark protocols.
@@ III-B Limitation of LLM’s Input Length
When inferring FSM from implementation using a pre-trained model, the complexity of the code and the sheer volume of files involved can hinder the model from producing accurate results. Protocol implementations typically comprise components such as state machine, configuration, encryption, etc. However, the code related to FSM is often concentrated in specific files. For instance, the message types are usually defined in a structure. Identifying these specific files can narrow the scope of the code.
Finding 2: The limitation of context window renders inferring state machine via LLM a challenging endeavor.
III-C LLM should be Fine-grained adjusted for FSM inference
Intuitively, augmented LLMs can deal with more specific tasks than pre-trained models, e.g., preprocessing the input, guiding the model with eticulously crafted prompts.
LLMPF [14] leverages the syntactic information generated by the LLM to guide protocol fuzzing. Rather than utilizing the LLM augmented with specific knowledge, LLMPF directly employs a pre-trained model. We compare the pre-trained model with the augmented model by employing a simple FSM inference case. For the pre-trained LLM, we adopt an approach akin to LLMPF, employing prompt engineering to guide the model in generating FSM. In contrast, we utilize a pre-trained large language model augmented with source code to infer the FSM. Initially, the pre-trained model is enhanced with the protocol implementation, followed by the application of the COT and the BAP to guide the model. The methodology for state machine inference using the augmented LLM is elaborated in Section IV. For the IKE protocol implementation, we selected strongSwan[23], a widely used solution in the industry.
TABLE III: The comparison between state machines inferred by pre-trained model and augmented model.
Models States Transitions Precision
Pre-trained LLM 4 12 43.48%
Augmented LLM 8 20 95.00%
Table III presents a comparison of the FSMs inferred by the pre-trained model and the augmented model. The detailed experimental methodology is discussed in Section V. The result shows that the augmented model outperforms the pre-trained model by 51.52% in terms of precision. The enhanced LLM demonstrates the capability to generate reliable state information. While the pre-trained LLM can offer some insights related to FSMs, its precision is limited to 43.48%. This discrepancy may be attributed to the hallucinations exhibited by LLMs when applied to domain-specific tasks.
Finding 3: Compared to the pre-trained model, the LLM augmented with source code exhibits a significantly enhanced capacity to generate a more reliable FSM.